OpenStudy (anonymous):

How would I graph y=2x-1 y= -1/2x+ 3/2

OpenStudy (anonymous):

vary x in any interval and collect the points of y and now write the points of x and y as (x,y) point this on graph paper and connect those points for ex take x=1 for that x y=1 so point is (1,1) and continue

OpenStudy (anonymous):

For y = 2x - 1: Make a table like: x | -2 -1 0 1 2 y | Try each x value in y = 2x - 1 to fill in the table. I'll start you off: y = 2(-2) -1 y = -4 -1 y = -5 Now the table is: x | -2 -1 0 1 2 y | -5 Repeat this a few times and you'll have a good idea of what the graph(line) is doing. (y = -1/2x + 3/2 can be done similarly)

OpenStudy (anonymous):

Alternatively, you might notice that each of those equations are in the form y = mx + b For example in y = 2x - 1: m = 2 and b = -1 and in y = -1/2x + 3/2: m = -1/2 and b = 3/2 Remember that m is the slope. Ex: when m=2, the graph goes up 2 for every 1 it goes to the right. Also, b is the y-intercept. Meaning that it crosses the y-axis at b. Ex: when b = -1, we know that the line crosses the y-axis at -1, which is the coordinate (0, -1)

OpenStudy (anonymous):

Sorry, that was a lot of text. Are you following?

OpenStudy (anonymous):

kinda..

OpenStudy (anonymous):

If you make a table, each set is a point on the graph. So you can connect all those points and have a good graph.